# Stability of Non-Hyperbolic Sets

### v1

@@ Abstract
In families of unfoldings we find codimension-one laminations consisting of maps exibiting an invariant, non-hyperbolic Cantor set containing a Collet–Eckmann point whose orbit is dense in the set. The topology of this Cantor set remains unchanged along the leaves of the lamination, ensuring codimension-one stability. Furthermore, within this lamination we identify a sublamination whose leaves contain maps with infinitely many sinks accumulating on the non-hyperbolic Cantor set containing the Collet–Eckmann point. These results provide a new perspective on the stability of chaotic non-hyperbolic dynamics in two-dimensional systems and reveal surprising coexistence phenomena that highlight the richness and complexity of such dynamics.
@@ 1 Introduction
This paper addresses two central questions in dynamical systems: the existence and stability of dynamically relevant sets. Given a dynamical system $(\mathcal{M},f)$ where $\mathcal{M}$ is a manifold and $f$ an evolution acting on it, one aims to identify the regions where the dynamics effectively take place—that is, the sets on which most orbits spend most of their time. Establishing the existence of such sets can already be a challenging task. Once their existence is established, one naturally seeks to understand whether these sets, together with their dynamical properties, persist under perturbations of the system. Are these sets isolated? Are their properties robust and observable across nearby systems? In other words, are they stable? Let us now make the notion of stability more precise.
@@
Given two dynamical systems $(M,f)$ and $(M,g)$ with relevant sets $A(f)$ and $A(g)$, we say that they are conjugate if there exists a homeomorphism $h:A(f)\to A(g)$ such that $h\circ f=g\circ h.$ One can think of this homeomorphism as a change of coordinates. For instance, if $A(f)$ and $A(g)$ are Cantor sets, the conjugacy map “rearranges” the pieces of one Cantor set to produce the other. Roughly speaking, this means that the two sets share the same topological structure, they are topologically equivalent.
@@
Within the space of dynamical systems, one can thus consider the class of maps whose relevant sets are topologically equivalent in this sense. The geometry of this class provides information about the stability of the corresponding sets. We say that the set $A(f)$ is codimension-$k$ stable if the class of systems conjugate to it forms a codimension-$k$ submanifold of the space of dynamical systems. This manifold is referred to as the manifold of stability.
@@
The sets exhibiting the strongest form of stability are sinks or closed attracting periodic orbits. Indeed, such sets are codimension-0 stable: their topology does not change under small perturbations of the system. More interesting examples arise in one-dimensional dynamics, such as $\mathcal{C}^{3}$ circle maps. In this class of systems, the topology of the relevant set (in this case, the entire circle) persists on codimension-one manifolds. The set is therefore codimension-$1$ stable. This phenomenon follows from one of the most profound results in one-dimensional dynamics, namely Herman’s Theorem, see [28]; for more modern expositions, see [20] and [30].
The stability of sets in one-dimensional dynamics is by now largely understood, and no major surprises are expected. Obtaining this complete picture required several decades of work and the efforts of many leading mathematicians. Although the main questions have been settled, a number of subtle and intriguing problems remain open.
In contrast, the situation in two dimensions is far richer and less complete. We now turn to the family of systems at the center of our study, namely the Hénon family (see Section 3 for a precise definition). In fact, in this paper we consider a broader class of families, of which the Hénon family is only one example. Specifically, we study unfoldings of maps with a homoclinic tangency and a transversal intersection. The precise definition is given in Subsection 2.1.
Which relevant sets are known to exist in such systems and what do we know about their stability?
@@
As expected, finitely many sinks can exist, and these are codimension-0 stable. Indeed, sinks are hyperbolic sets. The existence of infinitely many sinks has also been established (see [33]), together with some preliminary results concerning their stability (see [23]). More intricate dynamics have likewise been shown to occur. In particular, the existence of a period-doubling Cantor attractor—an invariant, non-hyperbolic, minimal set carrying a measure with zero Lyapunov exponent—has been proved (see [14, 31]). This set is known to be codimension- 1 stable. Furthermore, the existence of two period-doubling Cantor attractors has been demonstrated, and these are codimension-2 stable (see [36]).
@@
Even more complex invariant sets, known as strange attractors, have been shown to exist in this family. Strange attractors are invariant, non-hyperbolic, and chaotic sets (they carry a measure having positive Lyapunov exponent). The seminal work establishing their existence is [3], where a remarkably intricate and ingenious proof was developed. Similar constructions adapted to more general systems and further studies were later given in [7, 4, 6, 27, 32, 43, 44]. Coexistence of these sets is established in [5]. What makes these sets chaotic is the existence of a point satisfying the Collet–Eckmann condition, which, roughly speaking, asserts that the derivative of the map grows at least exponentially fast along the orbit of this point (see Definition 14 for a precise formulation).
A natural question then arises: what can be said about the stability of strange attractors? Their chaotic nature makes this question extremely challenging. It is known that, in parameter space, there exists a positive Lebesgue measure set of maps exhibiting such attractors; however, their topology changes as the parameter varies. So far, we lack the tools to clearly address this problem. The present paper aims to provide a partial answer to this question. A new construction and new tools are given. Furthermore, while trying to answer this question, we found new fascinating phenomena, namely the existence of maps with infinitely many sinks accumulating on a set containing a Collet-Eckmann point.
We now summarize our main results, which together describe the existence and stability of non-hyperbolic Cantor sets in families of unfoldings.”
@@
Let $\mathcal{M}$ be a two-dimensional manifold and let $F:\left([-t_{0},t_{0}]\times[-a_{0},a_{0}]\right)\times\mathcal{M}\to\mathcal{M}$ be a family of unfolding. There exists a codimension-$1$ lamination $\mathfrak{L}$ in parameter space such that, each leave $\ell\subset\mathfrak{L}$ and each map $f\in\ell$ have a set $A(f)$ satisfying the following.
@@
    The set $A(f)$ in an invariant, non-trivial, non-hyperbolic Cantor set.
@@
    The set $A(f)$ contains a point whose limit set coincides with the orbit of a transversal homoclinic intersection point.
@@
    Every leaves of $\mathfrak{L}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$. Moreover if $t\in[-t_{0},t_{0}]$, the closure of the set $L_{t}=\left\{(t,a)\in\mathfrak{L}\right\}$ is a Cantor set.
@@
    If $f$ and $\tilde{f}$ are in the same leave $\ell$ of the lamination $\mathfrak{L}$ then they are conjugate on their respective Cantor sets.
@@
Theorem $A$ is the summary of several results proved along the paper. In order: Lemma 22, Corollary 1, Lemma 24, Lemma 25, Theorem 6, Proposition 9, Theorem 4, Theorem 5.
Theorem A establishes the existence of an invariant, non-hyperbolic Cantor set that is codimension-1 stable. In particular, any two maps lying on the same leaf of the lamination are conjugate on their corresponding Cantor sets. In other words, along each leaf of the lamination, the corresponding Cantor sets are topologically conjugate, and their structure remains invariant.
These results are obtained through the introduction of dynamical partitions and the associated directed graphs. The inverse limit of these graphs is homeomorphic to the Cantor set. This combinatorial structure provides effective control over both the topological dynamics along the leaves and the measure-theoretical properties of the Cantor set. These inverse limits and directed graphs were introduced in [21]. In the present work, we adapt these ideas to the setting of families of maps exhibiting homoclinic tangencies.
To formally establish the codimension-one stability of the Cantor set, we introduce the notion of infinitely renormalizable maps (see Subsection 10.3). For simplicity, we refer to this process as renormalization, although it is, strictly speaking, a pre-renormalization. In particular, we study the first-return maps to the sets generating the dynamical partition, without performing any rescaling. Since our main focus lies on topological properties, this pre-renormalization suffices for our purposes. To the best of our knowledge, this represents the first attempt to formulate a generalized renormalization framework with two branches in the setting of two-dimensional systems. A formulation of renormalization for two-dimensional systems using only one branch has been explored in [8, 14, 16, 17, 36], but our construction differs in that it employs two branches and more closely resembles generalized renormalization as used for one-dimensional systems.
@@
Let us go back to analyze the statements in Theorem A. In particular, one should compare these results with those of [14], where the period-doubling Cantor attractor was shown to be codimension-1 stable. In that case, the topology of the Cantor set also persists along a codimension-1 leave; however, the constructed leave is very short. Numerical experiments suggest that the period-doubling locus should, in fact, form a long global curve, but proving this remains a challenging open problem. In contrast, our approach allows for the construction of global curves that are genuine manifolds of stability.
We next strengthen these results by showing the existence of a sub-lamination such that the corresponding non-hyperbolic invariant Cantor sets contain a Collet–Eckmann point with a dense orbit in the Cantor set. These sets, which exhibit chaotic dynamics, are likewise codimension-1 stable, i.e. their topology is preserved along the leaves of the lamination. This result represents an initial step toward the study of stability of strange attractors. To the best of our knowledge, no prior results are available in this direction. The proof of the theorem relies on the dynamical structure described above, together with a detailed analytical control along the tangency loci.
@@
There exists a sub-lamination $\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property. For each function $f\in\ell\subset\mathfrak{L}_{CE}$, the invariant non-hyperbolic Cantor set $A(f)$ contains a Collet-Eckmann point $c$ such that $\omega(c)=A(f)$. Moreover the closure $\overline{\mathfrak{L}_{CE}}=\overline{\mathfrak{L}}$.
Theorem B is proved in Theorem 7.
In pursuing the stability of chaotic non-hyperbolic sets, we encounter a striking coexistence phenomenon that further illustrates the richness of the dynamics in these systems. We prove the existence of a sub-lamination in which each leaf contains a map with infinitely many sinks of increasing period, accumulating on a non-hyperbolic Cantor set that contains a Collet–Eckmann point. This coexistence of infinitely many sinks with chaotic dynamics is both unusual and unexpected. The formal statement of the theorem is given below.
@@
There exists a sub-lamination $\mathfrak{L}_{N}\subset\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property. Each leave $\ell\subset\mathfrak{L}_{N}$ contains a map $f_{\ell}$ with infinitely many attracting periodic points, $p_{k}$ of arbitrarily high period, such that,
$\bigcap_{n}\overline{\bigcup_{k\geq n}\text{Orb}(p_{k})}=A(f_{\ell}).$
Moreover if $NC=\left\{f_{\ell}\right\}_{f_{\in}\ell\subset\mathfrak{L}_{N}}$, then the closure $\overline{{NC}}=\overline{\mathfrak{L}_{CE}}=\overline{\mathfrak{L}}$.
Theorem C is proved in Theorem 8. The proof relies on an intermediate result, Theorem 3, which establishes the existence of maps with infinitely many sinks in families of unfoldings. Maps exhibiting the Newhouse phenomenon have been constructed in various contexts. In particular, numerous examples of Baire [^1] A Baire set in a locally compact Hausdorff space X X is a countable intersection of open and dense subsets. sets of maps with infinitely many sinks in the space of dynamical systems are known; see, for instance, [1, 9, 13, 11, 15, 18, 19, 24, 25, 29, 33, 34, 41, 42].
@@
In the classical Newhouse construction, the coexistence of infinitely many sinks is produced by the persistence of homoclinic tangencies. The underlying mechanism relies on the intersection of stable and unstable Cantor sets whose thicknesses exceed one, ensuring that tangencies survive under perturbation.
Our approach is fundamentally different. To prove Theorem C, we avoid using the persistence of tangencies altogether. Instead, we introduce a quantitative condition (2.2) that involves only the eigenvalues of a saddle point. This spectral condition replaces the geometric thickness criterion of Newhouse and provides an explicit, verifiable rule for when the construction can be iterated. It is worth noting that, in the Hénon family, the set of maps satisfying the thickness condition and the set satisfying condition (2.2) intersect, but neither is contained in the other.
@@
The proof proceeds inductively: starting from a map with a sink and a homoclinic tangency, we show that the sink persists under small parameter changes, while the tangency gives rise to a new tangency and an additional sink. Iterating this mechanism yields infinitely many sinks accumulating on a non-hyperbolic Cantor set containing a Collet–Eckmann point.
The most significant difference between our approach and Newhouse’s original method [33] lies in the creation of the secondary tangency. In Newhouse’s construction, the persistence of tangencies guarantees that every nearby map possesses a tangency somewhere, but this dependence on the map is highly discontinuous, making the analysis delicate. Our method—inspired by [3, 5]—replaces this mechanism with the use of critical points and binding points, thereby avoiding any dependence on the persistence of tangencies.
Thus, our eigenvalue-based and constructive method offers an alternative route to the Newhouse phenomenon and clarifies how coexistence of chaos and stability can occur.

### v3

# arXiv:2511.12926v3[math.DS] 01 Mar 2026

@@ Abstract
In two-dimensional unfoldings of homoclinic tangencies, the parameter space contains codimension-1 laminations whose leaves consist of maps with invariant non-hyperbolic Cantor sets. These Cantor sets are wild both in the sense of Hofbauer–Keller and in the sense of Newhouse, and they contain Collet–Eckmann points with dense orbits. Hence, wildness and non-uniform chaotic hyperbolicity can coexist on a single invariant set, while persisting along codimension-1 manifolds.
In addition, each leaf of the lamination contains a map with infinitely many sinks accumulating on the Cantor set containing the Collet-Eckmann point. This surprising symbiosis of contraction and expansion may not, in fact, be pathological. Along the way, we introduce a generalized renormalization scheme for two-dimensional systems.
@@ 1 Introduction
This paper addresses two central questions in dynamical systems: the existence and stability of dynamically relevant sets. Given a dynamical system $(\mathcal{M},f)$ where $\mathcal{M}$ is a manifold and $f$ a map acting on it, one aims to identify the regions where the dynamics effectively take place, that is, invariant sets capturing the long-term behaviour of typical orbits. Establishing the existence of such sets can already be a challenging task. Once their existence is established, one naturally seeks to understand whether these sets, together with their dynamical properties, persist under perturbations of the system. Are these properties robust and observable across nearby systems? In other words, are they stable? We now introduce a notion of stability tailored to this setting.
@@
Given two dynamical systems $(\mathcal{M},f)$ and $(\mathcal{M},g)$ with relevant sets $A(f)$ and $A(g)$, we say that they are conjugate if there exists a homeomorphism $h:A(f)\to A(g)$ such that $h\circ f=g\circ h.$ One can think of this homeomorphism as a change of coordinates. For instance, if $A(f)$ and $A(g)$ are Cantor sets, the conjugacy map “rearranges” the pieces of one Cantor set to produce the other. In this sense, the two sets share the same topological structure and are therefore topologically equivalent.
@@
Within the space of dynamical systems, one can therefore consider the class of maps whose relevant sets are topologically equivalent. The structure of this class encodes the stability of the corresponding sets. We say that the set $A(f)$ is codimension-$k$ stable if the set of systems whose relevant sets are conjugate forms a codimension-$k$ submanifold of the space of dynamical systems. This submanifold is called the manifold of stability.
The sets exhibiting the strongest form of stability are closed attracting periodic orbits and, more generally, hyperbolic invariant sets. Indeed, such sets are codimension-0 stable, meaning that their topology does not change under small perturbations of the system.
@@
More subtle phenomena already arise in one-dimensional dynamics. For instance, in the class of $\mathcal{C}^{3}$ circle maps, the relevant invariant set is the entire circle, but its topological structure is no longer stable under arbitrary perturbations. Instead, it persists only along codimension-1 manifolds in parameter space, and is therefore codimension-1 stable. This phenomenon follows from one of the most profound results in one-dimensional dynamics, namely Herman’s Theorem, see [40], and for more modern expositions [26, 43].
The stability of invariant sets in one-dimensional dynamics is by now largely understood, and no fundamentally new phenomena are expected. That said, a number of subtle and interesting questions remain open, mostly concerning finer properties of the dynamics and their dependence on parameters.
In contrast, the situation in two dimensions is much richer and far from completely understood. Several mechanisms of instability may interact in nontrivial ways, making questions of existence and stability more delicate. We now turn to the family of systems at the center of our study, namely unfoldings of maps with a homoclinic tangency and a transversal intersection. Such unfoldings provide a natural setting in which complex invariant sets and coexistence phenomena arise. The precise definition is given in Subsection 2.1. A central example is provided by the Hénon family (see Section 3 for a precise definition). What relevant invariant sets arise in this context, and what is known about their stability?
As expected, a finite number of attracting periodic orbits (sinks) may coexist, and these are codimension-0 stable. Indeed, sinks are hyperbolic sets. More surprisingly, the existence of infinitely many sinks has also been established (see [48]), together with initial results concerning their stability (see [35]).
@@
Beyond periodic behaviour, non periodic dynamics have likewise been shown to occur. In particular, the existence of a period-doubling Cantor attractor, an invariant, non-hyperbolic, minimal set carrying a measure with zero Lyapunov exponent, has been proved (see [17, 33, 45]). This set is known to be codimension-1 stable. In addition, maps exhibiting two period-doubling Cantor attractors have been constructed, and these are codimension-2 stable (see [51]).
@@
Even more intricate invariant sets, known as strange attractors, have been shown to exist in Hénon-like families. Strange attractors are invariant, non-hyperbolic, and chaotic sets, in the sense that they support an invariant measure with positive Lyapunov exponent. Their existence was first established in the seminal work [4], which introduced a new and highly nontrivial construction. Similar constructions and further developments were later given in [5, 7, 8, 39, 47, 58, 59], and the coexistence of multiple strange attractors was established in [6].
In many constructions, the analysis of chaotic dynamics relies on the identification of a point satisfying the Collet–Eckmann condition. Roughly speaking, this condition asserts that the derivative of the map grows exponentially fast along the orbit of this point (see Definition 12.0.1 for a precise formulation).
A natural question then arises: what can be said about the stability of strange attractors? Because of their chaotic nature, this question is particularly challenging. It is known that, in parameter space, there exists a positive Lebesgue measure set of maps exhibiting such attractors. However, the structure of these attractors typically changes as the parameter varies. At present, we lack general tools to determine whether, and in what sense, such chaotic sets can be stable. The present paper provides a first answer to this question for a class of non-hyperbolic chaotic invariant sets. Our approach is to identify a precise notion of stability for dynamics that lie well beyond the uniformly hyperbolic setting.
Before summarizing our main results, we explain the viewpoint underlying this work. The invariant sets we construct display several types of dynamical behaviour that are commonly regarded as wild. First, they are statistically wild in the sense of Hofbauer and Keller, as stated in Theorem A-(iv), in that a generic orbit in the set is dense, while from a statistical point of view the behavior is degenerate, meaning that typical orbits accumulate on a very singular measure, as described for instance in [15, 41, 44]. Second, they are wild in the sense of Newhouse, see Theorem A-(viii), in the classical terminology of [12, 48, 50]: arbitrarily small perturbations exhibit infinitely many sinks. Moreover, all constructed Cantor sets contain a Collet–Eckmann point with dense orbit, see Theorem B. Furthermore, we show that these sets persist along the leaves of a codimension-1 lamination. In summary, wild and non-hyperbolic chaotic behavior can coexist on a single invariant set while persisting along codimension-1 manifolds.
In addition, Theorem C states that each leaf of the lamination contains a map with infinitely many sinks accumulating on the Cantor set containing the Collet-Eckmann point. This surprising symbiosis of contraction and expansion might not even be pathological, and motivates the Conjecture stated below.
The main contribution of this paper is to show that all these phenomena can coexist on a single invariant Cantor set and, moreover, persist along codimension-one laminations in parameter space. Theorems A, B, and C make this precise by progressively incorporating these behaviours and establishing a unified notion of stability for such highly non-hyperbolic dynamics.
We now summarize the main results of the paper.
@@
Let $\mathcal{M}$ be a two-dimensional manifold and let $F:\left([-t_{0},t_{0}]\times[-a_{0},a_{0}]\right)\times\mathcal{M}\to\mathcal{M}$ be an unfolding with $t_{0},t_{0}]\times\left\{0\right\}$ as homoclinic tangency locus of a saddle point. There exists a codimension-$1$ lamination $\mathfrak{L}$ in parameter space such that, each leaf $\ell\subset\mathfrak{L}$ and each map $f\in\ell$ have a set $A(f)$ satisfying the following.
@@
    The set $A(f)$ in an invariant, non-hyperbolic Cantor set.
@@
    The set $A(f)$ contains a saddle point together with a corresponding transversal homoclinic intersection.
@@
    Every leaf of $\mathfrak{L}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\to[-a_{0},a_{0}]$. Moreover if $t\in[-t_{0},t_{0}]$, the closure of the set $L_{t}=\left\{(t,a)\in\mathfrak{L}\right\}$ is a Cantor set.
@@
    If $f$ and $\tilde{f}$ are in the same leaf $\ell$ of the lamination $\mathfrak{L}$ then they are conjugate on their respective Cantor sets $A(f)$ and A(f~)A($\tilde${f}).
@@
  • (viii)
    Every open neighborhood of $f\in\mathfrak{L}$ contains maps with infinitely many sinks.
@@
Theorem A is the summary of several results proved in the paper. In order: Lemma 9.3.1, Corollary 11.0.2, Lemma 9.3.4, Lemma 9.3.5, Theorem 11.0.1, Proposition 9.1.3, Theorem 10.2.3, Theorem 10.3.3, Remark 9.1.5.
Theorem A establishes the existence of an invariant, non-hyperbolic Cantor set that is codimension-1 stable. In particular, any two maps lying on the same leaf of the lamination are conjugate on their corresponding Cantor sets, so that the dynamical structure is preserved along the leaf. This stability is striking given the wildness of these Cantor sets. Indeed, they are Newhouse-unstable and support only a trivial invariant probability measure, namely the Dirac mass at a saddle point. In this sense, Theorem A already reveals a first form of stability for dynamics with wild properties.
@@
These results should be compared with those of [17], where the period-doubling Cantor attractor was shown to be codimension-1 stable. In that case, the topology of the Cantor set likewise persists along a codimension-1 leaf; however, the constructed leaf is very short. Numerical experiments suggest that the period-doubling locus should in fact form a long global curve, but establishing this rigorously remains a challenging open problem. By contrast, our approach exploits the unstable expansion at the saddle point to construct a lamination in parameter space whose leaves have uniform size and act as manifolds of stability.
The underlying idea behind the proof of Theorem A is a renormalization scheme, which can be understood as an iterative procedure based on induced maps. More precisely, one studies the dynamics by restricting the map to carefully chosen regions of phase space and considering the map induced by the first return to these regions. These induced maps describe the behaviour of the system at progressively smaller scales. Renormalization is defined by iterating this construction: starting from the original map, one passes to an induced map on a suitable domain, rescales it, and then repeats the same induced-map procedure. This iteration thus defines a renormalization operator acting on the space of maps.
The regions selected at successive steps form a nested sequence that progressively refines the phase space. This nested structure gives rise to what are known as dynamical partitions, which encode the combinatorial and geometric organization of the dynamics across scales.
In our study, we focus primarily on the analysis of these dynamical partitions and rely less on the associated renormalization operator itself. The combinatorics of our renormalization scheme is intrinsically unbounded. In such unbounded settings, a central difficulty is to control the geometry of the partitions and to obtain uniform estimates on how the relevant quantities vary with the parameters.
This difficulty is overcome in Proposition 4.2.2 and its proof. Roughly speaking, this result provides quantitative control on how certain distinguished points (critical values) move with the parameters, and how the corresponding dynamical partitions are shaped. While the proof of Proposition 4.2.2 is very technical, it contains the core of our analysis.
The combinatorics of the renormalization scheme and the ideas underlying Proposition 4.2.2 are inspired by [4]. A fundamental difference, however, is that we only need to control a linearly growing number of critical points, whereas in [4] an exponentially growing number of such points must be handled. This reduced combinatorial complexity allows the dynamics to be organized in a more rigid and transparent way, which enables the constructions described below.
The lamination described in Theorem A consists of infinitely renormalizable maps (see Subsection 10.3 for a formal definition). As explained above, the renormalization procedure associated with such maps produces a sequence of dynamical partitions at finer and finer scales. The intersection of these nested partitions is an invariant Cantor set. Rather than studying the dynamics on the invariant Cantor set directly, we encode it using a combinatorial model based on the dynamical partitions. To this end, we associate a directed graph to each level of the renormalization procedure. The vertices of the graph correspond to the elements of the dynamical partition at that level, and the edges record how these elements are mapped onto one another by the dynamics. In this way, each graph provides a finite combinatorial description of the dynamics at a given scale.
The inverse limit of these graphs carries a natural dynamics and is a Cantor set. This dynamical system is conjugate to the dynamics induced by the original map on its invariant Cantor set. As a result, this combinatorial description allows effective control of the topological, geometric, and measure-theoretical properties of the Cantor sets. Sequences of directed graphs of this type were introduced in [32]. In the present work, we extend this construction to the setting of families of maps exhibiting homoclinic tangencies.
We would like to stress at this point that our approach represents the first successful attempt to formulate a generalized renormalization framework with two branches in the setting of two-dimensional systems with positive entropy. Renormalization schemes for two-dimensional systems using only one branch have been explored in [1, 9, 17, 20, 21, 22, 23, 27, 28, 29, 30, 31, 33, 45, 46, 51]. Our construction differs in that it employs two branches and, in this respect, is closer in spirit to the generalized renormalization frameworks developed for one-dimensional systems.
We next extend the list of properties of the Cantor sets described in Theorem A. In particular, we show that these non-hyperbolic invariant Cantor sets exhibit chaotic dynamics, in the sense that they contain a Collet–Eckmann point with dense orbit. Moreover, these sets remain codimension-1 stable, and their Collet–Eckmann points persist continuously along the leaves of the lamination. This shows that strong chaotic behaviour, manifested by exponential growth along a dense orbit, is compatible with a robust topological structure in parameter space. Theorem B thus represents a first step toward the study of the stability of chaotic non-hyperbolic invariant sets. At the present stage, however, since our analysis controls only a linearly growing number of critical points, it would be premature to formulate a general conjecture concerning the stability of strange attractors.
@@
There exists a sub-lamination $\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property. For each leaf $\ell\subset\mathfrak{L}_{CE}$ and for each map $f\in\ell$, the invariant non-hyperbolic Cantor set $A(f)$ contains a Collet-Eckmann point $c_{f}$ satisfying the following properties.
  • The omega-limit set $\omega(c_{f})=A(f)$.
  • The point $c_{f}$ depends continuously on $f\in\mathfrak{L}_{CE}$.
[…]
Moreover, every leaf of $\mathfrak{L}_{CE}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$ and if $t\in[-t_{0},t_{0}]$, the closure of the set $L^{CE}_{t}=\left\{(t,a)\in\mathfrak{L}_{CE}\right\}$ is a Cantor set.
Theorem B is a reformulation of Theorem 12.0.2. In pursuing the stability of chaotic non-hyperbolic sets, we encounter a striking coexistence phenomenon that further illustrates the richness of the dynamics in these systems. We prove that each leaf of the lamination contains a map with infinitely many sinks of increasing period. Moreover these periodic sinks accumulate on the non-hyperbolic Cantor set that contains a Collet–Eckmann point. The coexistence of infinitely many sinks with chaotic dynamics is surprising. The formal statement of the theorem is given below.
@@
There exists a sub-lamination $\mathfrak{L}_{N}\subset\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property. Each leaf $\ell\subset\mathfrak{L}_{N}$ contains a map $f_{\ell}$ with infinitely many attracting periodic points $p_{g}$, of arbitrarily high period, such that,
$\bigcap_{n}\overline{\bigcup_{g\geq n}\text{Orb}(p_{g})}=A(f_{\ell}).$
Moreover, every leaf of $\mathfrak{L}_{N}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$ and if $t\in[-t_{0},t_{0}]$, the closure of the set $L^{N}_{t}=\left\{(t,a)\in\mathfrak{L}_{N}\right\}$ is a Cantor set.
Theorem C is a reformulation of Theorem 12.0.3.
For a map with infinitely many sinks, it is natural to expect that each sink arises from a local return map which, after rescaling, is Hénon-like. The creation of further sinks at smaller scales therefore requires that these return maps be parameter unstable, as stability under parameter variations would prevent the appearance of new sinks. Since parameter instability is often associated with dynamical instability of the underlying system, this leads to the following conjecture.
Conjecture.
Let $f$ be a diffeomorphism of a compact two-dimensional manifold with infinitely many attracting periodic points $p_{g}$, of arbitrarily high period. If the set
$A=\bigcap_{n}\overline{\bigcup_{g\geq n}\text{Orb}(p_{g})},$
is transitive then it contains a Collet-Eckmann point with dense orbit.
The proof of Theorem C relies on an intermediate result, Theorem 6.2.7, which establishes the existence of maps with infinitely many sinks in unfoldings. The coexistence of infinitely many attracting periodic orbits is known as the Newhouse phenomenon, first discovered by Newhouse in [48]. Since its discovery, the Newhouse phenomenon has been observed in a wide range of settings. In particular, many constructions yield Baire sets of maps with infinitely many sinks in the space of dynamical systems; see, for instance, [2, 10, 13, 16, 18, 24, 25, 36, 37, 42, 48, 49, 56, 57]. These results demonstrate that the phenomenon is abundant from a topological point of view, although its detailed dynamical structure often remains poorly understood.
@@
In Newhouse’s original construction, the coexistence of infinitely many sinks is produced by the persistence of homoclinic tangencies. The key mechanism relies on the interaction of stable and unstable Cantor sets whose thicknesses exceed one, which guarantees that tangencies persist under perturbations of the system. This persistence, in turn, allows the repeated creation of sinks.
@@
Our proof of Theorem C follows a different, more explicit strategy. The argument proceeds inductively. Starting from a map with a sink and a homoclinic tangency, we show that the sink persists under small parameter variations, while the tangency gives rise to a new, explicitly constructed tangency, called secondary tangency, and an additional sink. The construction of the secondary tangency is intrinsically related to the renormalization scheme. Repeating this procedure produces infinitely many sinks of increasing period, which accumulate on a non-hyperbolic Cantor set containing a Collet–Eckmann point.
The main difference between our approach and Newhouse’s original method [48] lies precisely in the treatment of the secondary tangency. In the classical construction, the persistence of tangencies guarantees that nearby maps exhibit tangencies somewhere, but the location of these tangencies depends highly discontinuously on the map, making a detailed analysis difficult. By contrast, our method, inspired by [4, 6], replaces persistence arguments with an explicit construction of the next tangency at each step. This constructive approach provides an alternative route to the Newhouse phenomenon and clarifies how the coexistence of chaotic dynamics and stable periodic behaviour can arise within the same system. Moreover, the explicit nature of the construction allows for a precise geometric analysis in both parameter space and phase space. To the best of our knowledge, this is the first time that the asymptotic dynamics of Newhouse sinks accumulating on a chaotic non-hyperbolic set is described in such detail.